OpenStudy (anonymous):

Slope = -2 (3,5) is on the line Find the y intercept :y=mx+b 5=(-2) ( _______) +b 5=______+b +______+______ ______+b

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@hockeychick23

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

so you just want the y intercept or ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

like i just need to finish the equation

OpenStudy (dtan5457):

ok well then in theory y=mx+b right? so the y here would be the coordinate 5 5=-2+b solving for b it's 7 so the equation would be 5=-2x+7 which is the same as y=-2x+7

OpenStudy (dtan5457):

y intercept = 7